Need help
The Young's modulus for a femur is 9 times 10^9 N/
Answer #1

Youngs Modulus (E) is related to forces (F) as

E =FL/AdeltaL

A= cross sectional area = 8 cm2= 8*10-4 m2, L= initial length= 43cm =43/100 =0.43cm F= mg/2 ( half the weight of the body) , m= mass of body ( asumed as 70kg) , F= 70*9.8/2 =343 N

deltaL= change in length = FL/AE = 343*0.43/(9*109*8*10-4)=2.048*10-5 m

F= half the weight of